Prove that any even power of a Nonzero integer is positive. That is, if X does not equal to zero in integers, then X2n is positive for every positive integer n.


1
Expert's answer
2021-05-14T11:07:06-0400

"x \neq 0, x^2>0" because

"\ if\ x>0, x^2=x*x>0,\\if\ x<0, x^2=x*x=(-x)*(-x)>0"

"x^{2n}=(x^2)^n>0"

since raising a positive number to a power keeps it positive